# InfiniBench: Infinite Benchmarking for Visual Spatial Reasoning with Customizable Scene Complexity

### v1

@@ Abstract
Modern vision-language models (VLMs) are expected to have abilities of spatial reasoning with diverse scene complexities, but evaluating such abilities is difficult due to the lack of benchmarks that are not only diverse and scalable but also fully customizable. Existing benchmarks offer limited customizability over the scene complexity and are incapable of isolating and analyzing specific VLM failure modes under distinct spatial conditions. To address this gap, instead of individually presenting benchmarks for different scene complexities, in this paper we present InfiniBench, a fully automated, customizable and user-friendly benchmark generator that can synthesize a theoretically infinite variety of 3D scenes with parameterized control on scene complexity. InfiniBench uniquely translates scene descriptions in natural language into photo-realistic videos with complex and physically plausible 3D layouts. This is achieved through three key innovations: 1) a LLM-based agentic framework that iteratively refines procedural scene constraints from scene descriptions; 2) a flexible cluster-based layout optimizer that generates dense and cluttered scenes previously intractable for procedural methods; and 3) a task-aware camera trajectory optimization method that renders scenes into videos with full object coverage as VLM input. Experiments demonstrate that InfiniBench outperforms state-of-the-art procedural and LLM-based 3D generation methods in prompt fidelity and physical plausibility, especially in high-complexity scenarios. We further showcased the usefulness of InfiniBench, by generating benchmarks for representative spatial reasoning tasks including measurement, perspective-taking and spatiotemporal tracking.
@@ 1 Introduction
Visual spatial reasoning yang2025thinking, which comprehends the placement, orientation, and relationships of objects within a 3D space, is essential for AI to understand the real-world scenes from visual inputs such as images and video streams bai2023qwen; chen2024internvl; hurst2024gpt. Unlike vision tasks on semantic perception of objects (e.g. image classification devadason2025comprehensive, segmentation csurka2022semantic; barbosa2023threefoldreviewdeepsemantic and captioning ghandi2023deep; hu2022scaling), visual spatial reasoning requires abilities in geometric cognition, including spatial relation understanding, mental rotation and perspective taking stogiannidis2025mind.
@@
However, current vision-language models (VLMs) struggle to bridge the gap between 2D visual perception and 3D spatial reasoning, especially in scenes with high complexity chen2025gleam. Such scene complexity is a multi-faceted concept that describes how the difficulty of understanding the scene’s spatial arrangements varies, and it includes three dimensions: 1) Compositional complexity about the scene content such as the number and variety of objects, which could result in visual clutter chen2025gleam; 2) Relational complexity about the intricacy of spatial relationships between objects, as more complicated spatial arrangement or higher ratio of scene occupancy challenges the VLM’s capacity of accurately grounding the spatial language chen2025spatial; 3) Observational complexity about how the scene is presented, encompassing challenges introduced by extreme viewpoints and object occlusion yu2025far. These three dimensions disentangle different targets in human spatial perception, including the scene elements (compositional), the interconnections between these elements (relational), and the viewpoint-dependent nature of perception (observational) greene2023scene.
@@
To thoroughly understand the VLM’s abilities of spatial reasoning in these dimensions is a big challenge, because they need benchmarks that are not only diverse and scalable but also fully customizable. Benchmarks should enable full isolation and flexible manipulation of scene properties, such as object density, relative positioning and levels of occlusion levels, through a set of configurable parameters within a large combinatorial space. For example, to evaluate VLM’s robustness to irrelevant objects, the benchmark should be able to systematically increase the number of irrelevant objects, while retaining the task-related objects unchanged. This fine-grained customizabilility will allow us moving beyond the aggregate performance metrics (e.g., average task accuracy yang2025thinking), to uncover the specific failure modes of VLMs under distinct spatial conditions.
@@
Existing benchmarks face significant limitations in customizablility, scalability, and semantic richness. Real-world datasets yin2023lamm; hong20233d; jia2025omnispatial; wu2025spatial; chen2024spatialvlm; jia2025omnispatial, although photorealistic, are costly to scale and lack parameterized control. A better alternative is to use synthetic data. Early methods use 3D rendering engines such as Blender blender and IsaacSim isaacSim to procedurally generate 3D scenes, but lack realism and fail to capture the real-world functional and aesthetic constraints wang20233d. Recent techniques address these limitations by using 3D-aware diffusion models meng2025lt3sd; liu2024pyramid; go2025videorfsplat; go2025splatflow; engstler2025syncity; li2025worldgrow; xiang2025structured with text yang2025prometheus; go2025videorfsplat or image priors engstler2025syncity; xia2025scenepainter for visual richness, but lack semantic annotations to generate QA pairs for reasoning tasks and fail to guarantee physical scene coherence. LLM-based scene generation ran2025direct; bucher2025respace; sun2025layoutvlm; zhang2025scene can provide better semantic grounding, but falter on generating more complex scenes with correct layouts, due to LLMs’ inherent limitations in spatial reasoning. Optimization-based procedural frameworks, such as Infinigen raistrick2024infinigen and ProcTHOR deitke2022, similarly fail to generate complex and cluttered scenes, and depend on professional tuning in practical use.
@@
To achieve this, we first developed a new LLM-based agentic framework that wraps an advanced optimizer to generate and refine the 3D scene constraints, according to the given prompt of scene descriptions. The refined scene constraints are then transformed into 3D scenes by a custom optimization engine, which iteratively samples and places objects to satisfy all the semantic and physical rules in the scene constraints. In particular, we replace the conventional method of hierarchical optimization xu2002constraint; yu2011make; merrell2011interactive with a more flexible cluster-based optimization strategy, to allow creating significantly more complex and dense scenes that were previously intractable. In this way, InfiniBench allows broad but effortless customization of the generated 3D scene from multiple perspectives (e.g., number and type of objects, spatial relationships, room layout, etc), without requiring professional knowledge or complicated tuning.
Another gap between 3D scene generation and VLM benchmarking is that outputs of current scene generation methods, typically 3D models in Blender files or point clouds, are not directly consumable by VLMs. To generate image sequences or video streams that VLMs can properly perceive, we developed a new method of camera trajectory optimization, ensuring that all task-relevant objects and their spatial relationship are clearly visible and covered in camera views. Following the principle of frontier-based exploration yamauchi1998frontier; lluvia2021active; yamauchi1997frontier, our algorithm incrementally builds the camera path by addressing each task-relevant object in sequence. For each object, it samples and evaluates multiple viewpoints and selects the best one that offers an unobstructed view after passing an occlusion check. A navigable path is then planned to link these viewpoints.
@@
We further showed that customizable benchmarks generated by InfiniBench facilitate in-depth analysis of VLM’s failure modes in various spatial reasoning tasks, including measurement (e.g. metric distance estimation), perspective-taking (e.g. object counting) and spatiotemporal tracking (e.g. identification of order identification) brown2025sims. Such customizability in benchmarking would also be a key enabler of enhancing VLM’s spatial reasoning capabilities, via spatially grounded training or prompt tuning.
@@ 2.1 Variability of Scene Complexity
Existing spatial reasoning benchmarks chen2024spatialvlm; wu2025spatial; jia2025omnispatial often categorize their data samples merely by task type or define scene complexity in simple terms, such as the number of rooms chen2025gleam. Such coarse-grained variability conflates multiple dimensions of the scene complexity, and hence obscures the precise reasons of VLM failures in reasoning tasks, which may include an excessive number of distractors (compositional complexity), an unusually intricate object arrangement (relational complexity), or heavily occluded viewpoints (observational complexity). Instead, in InfiniBench, we allowed much more fine-grained variability of scene complexity in these dimensions, by using separate sets of scene parameters (specified by users in natural language) to explicitly represent complexity levels of these dimensions and generating scenes accordingly.
@@ 2.2 Layout Generation in 3D Scene Generation
The main focus of InfiniBench design is the creation of scene layout, which defines the placement, orientation, and relationships of objects in the scene. One can leverage the reasoning capabilities of LLMs to directly produce a layout and then instruct a 3D engine to place assets in the scene ran2025direct; bucher2025respace; sun2025layoutvlm; zhang2025scene. To improve the coherence of generated layouts, some works guide LLMs with structural templates feng2023layoutgpt or represent scenes as graphs to facilitate compositional generation ccelen2024design; deng2025global; gao2024graphdreamer. However, the generated scene’s complexity is constrained by current LLMs’ flawed spatial logic. As shown in Figure 1, as the number of objects and the intricacy of constraints increase, these methods frequently produce physically implausible layouts, such as those with unrealistic object orientations, out-of-boundary objects or overlapped objects.
To address this challenge and ensure the scalability of generating complex scenes, we proposed a more robust approach that separates high-level planning from low-level execution. Instead of tasking the LLM to generate the exact scene layout, we use LLM to generate high-level constraints about object arrangement. These constraints are then fed into a dedicated optimizer that generates a physically coherent 3D scene. By using optimization-based procedural generation frameworks raistrick2024infinigen; deitke2022 that are effective in 3D scene synthesis zhao2021luminous; yu2011make; yu2015clutterpalette, we designed an improved optimizer for generating complex scenes and wrap this optimizer within an LLM-based agentic framework, making the entire pipeline both powerful and user-friendly.
@@ 2.3 Camera Trajectory Generation
Optimization-based methods sample numerous camera parameters with respect to a predefined objective, to maximize the object visibility bares2000virtual; halper2001camera. While offering fine-grained control, this approach is expensive and struggles with defining an objective function that robustly captures the visual requirements for complex spatial reasoning tasks. Another alternative is to train a generative AI model courant2024exceptional; jiang2021camera; wang2024motionctrl. However, existing datasets for such training are primarily for aesthetically pleasing and object-centric photo shots, e.g., smooth orbits around a single object, rather than trajectories with controlled coverage of multiple objects. Some other methods xie2023gait; dehghanian2025camera used reinforcement learning to train an agent for scene navigation without large training data, but needs to retrain the agent for each new scene.
Instead, our approach is inspired by frontier-based exploration yamauchi1997frontier; yamauchi1998frontier; lluvia2021active, a lightweight technique widely used in robotic navigation and mapping. In InfiniBench, we redefine frontier as object clusters, and sample viewpoints targeting these frontiers accordingly to strategically manage factors such as object occlusion.
@@ 3.1 Agentic Generation of Scene Constraints
A primary barrier to traditional procedural generation of 3D scenes is the need of manually scripting complex scene constraints, a process that requires significant technical expertise, especially when the scene complexity grows. InfiniBench democratizes this process, as shown in Figure 3, by using a LLM agent to translate scene descriptions in natural language into procedural constraints. To ensure that the generated constraints are machine-readable, we teach the LLM agent with domain knowledge of 1) a comprehensive list of available procedural APIs and 2) a set of few-shot examples. The information about procedural APIs covers detailed syntax and functionality with respect to the procedural 3D generation framework being used (e.g., Infinigen raistrick2024infinigen and ProcTHOR deitke2022). The few-shot examples demonstrate the translation from high-level user requests (e.g., “a cluttered office desk”) to constraint definitions (e.g., set_object_count(monitor, 3), on_top_of(keyboard, desk)). More details about such in-context learning are in Appendix A.
@@ 3.2 Layout Optimization for Complex Scenes
When the scene complexity grows, hierarchical optimization methods used in the existing procedural generation frameworks such as Infinigen raistrick2024infinigen and ProcTHOR deitke2022 would fail, because they first fix the positions of large objects (e.g. desks and beds) and usually result in unsolvable states where there is no space for smaller objects (e.g. chairs around the desk), as shown in Figure 5(a).
@@ 3.3 Camera Trajectory Optimization
Given a generated 3D scene and a spatial reasoning task, the objective of our camera trajectory optimization in InfiniBench is to generate the shortest possible camera path that provides a clear, full and unobstructed view of every task-relevant object. Our method is inspired by frontier-based exploration yamauchi1997frontier; yamauchi1998frontier; lluvia2021active, a classical robotic navigation technique. Unlike traditional approaches that explore frontiers using occupancy grids, we define “frontiers” as the set of unvisited target objects. We then develop a viewpoint sampling method to find optimal views for these targets while avoiding occlusion. As shown in Figure 6, this process consists of the following steps:
@@
    Path planning: A collision-free path is computed from the current camera position to the selected viewpoint using Dijkstra’s algorithm in 3D space.
@@
I-Design ccelen2024design 0.95 28.4 0.72 1.6 0.4 0.92 27.9 0.68 3.4 5.1 0.90 27.1 0.61 6.9 10.3
Holodeck yang2024holodeck 0.97 30.1 0.81 1.9 2.1 0.93 29.5 0.77 4.6 3.1 0.88 28.8 0.71 7.7 9.4
LayoutGPT feng2023layoutgpt 0.96 29.8 0.79 1.3 2.4 0.95 29.1 0.74 3.7 4.6 0.93 28.3 0.67 4.5 13.5
Luminous zhao2021luminous 0.87 27.5 0.75 0.0 0.0 0.84 26.8 0.70 0.0 0.1 0.42 26.2 0.63 0.0 0.0
Infinigen raistrick2024infinigen 0.92 31.0 0.88 0.0 0.1 0.87 30.4 0.84 0.0 0.2 0.64 29.7 0.79 0.0 0.2
@@
I-Design ccelen2024design 0.90 28.6 0.73 0.1 0.4 0.87 28.0 0.67 1.3 3.1 0.79 27.2 0.59 5.3 6.4
Holodeck yang2024holodeck 0.78 30.4 0.82 0.3 0.2 0.80 29.7 0.76 0.9 5.4 0.75 28.9 0.69 3.1 8.7
LayoutGPT feng2023layoutgpt 0.81 29.9 0.80 0.9 0.2 0.83 29.3 0.73 1.7 6.8 0.85 28.5 0.65 5.6 9.6
Luminous zhao2021luminous 0.74 27.8 0.76 0.0 0.0 0.64 27.1 0.69 0.0 0.0 0.43 26.4 0.61 0.0 0.0
Infinigen raistrick2024infinigen 0.83 31.2 0.89 0.0 0.0 0.71 30.6 0.85 0.2 0.1 0.49 29.9 0.78 0.1 0.2

### v2

@@ Abstract
Modern vision-language models (VLMs) are expected to have abilities of spatial reasoning with diverse scene complexities, but evaluating such abilities is difficult due to the lack of benchmarks that are not only diverse and scalable but also fully customizable. Existing benchmarks offer limited customizability over the scene complexity and are incapable of isolating and analyzing specific VLM failure modes under distinct spatial conditions. To address this gap, instead of individually presenting benchmarks for different scene complexities, in this paper we present InfiniBench, a fully automated, customizable and user-friendly benchmark generator that can synthesize a theoretically infinite variety of 3D scenes with parameterized control on scene complexity. InfiniBench uniquely translates scene descriptions in natural language into photo-realistic videos with complex and physically plausible 3D layouts. This is achieved through three key innovations: 1) a LLM-based agentic framework that iteratively refines procedural scene constraints from scene descriptions; 2) a flexible cluster-based layout optimizer that generates dense and cluttered scenes previously intractable for procedural methods; and 3) a task-aware camera trajectory optimization method that renders scenes into videos with full object coverage as VLM input. Experiments demonstrate that InfiniBench outperforms state-of-the-art procedural and LLM-based 3D generation methods in prompt fidelity and physical plausibility, especially in high-complexity scenarios. We further showcased the usefulness of InfiniBench, by generating benchmarks for representative spatial reasoning tasks including measurement, perspective-taking and spatiotemporal tracking. The source codes of InfiniBench are available at https://github.com/pittisl/infinibench, and some sample benchmarks being generated by InfiniBench are available at https://huggingface.co/datasets/Haoming645/infinibench.
@@ 1 Introduction
Visual spatial reasoning [yang2025thinking], which comprehends the placement, orientation, and relationships of objects within a 3D space, is essential for AI to understand the real-world scenes from visual inputs such as images and video streams [bai2023qwen, chen2024internvl, hurst2024gpt]. Unlike vision tasks on semantic perception of objects (e.g. image classification [devadason2025comprehensive], segmentation [csurka2022semantic, barbosa2023threefoldreviewdeepsemantic] and captioning [ghandi2023deep, hu2022scaling]), visual spatial reasoning requires abilities in geometric cognition, including spatial relation understanding, mental rotation and perspective taking [stogiannidis2025mind].
@@
However, current vision-language models (VLMs) struggle to bridge the gap between 2D visual perception and 3D spatial reasoning, especially in scenes with high complexity [chen2025gleam]. Such scene complexity is a multi-faceted concept that describes how the difficulty of understanding the scene’s spatial arrangements varies, and it includes three dimensions: 1) Compositional complexity about the scene content such as the number and variety of objects, which could result in visual clutter [chen2025gleam]; 2) Relational complexity about the intricacy of spatial relationships between objects, as more complicated spatial arrangement or higher ratio of scene occupancy challenges the VLM’s capacity of accurately grounding the spatial language [chen2025spatial]; 3) Observational complexity about how the scene is presented, encompassing challenges introduced by extreme viewpoints and object occlusion [yu2025far]. These three dimensions disentangle different targets in human spatial perception, including the scene elements (compositional), the interconnections between these elements (relational), and the viewpoint-dependent nature of perception (observational) [greene2023scene].
@@
To thoroughly understand the VLM’s abilities of spatial reasoning in these dimensions is a big challenge, because they need benchmarks that are not only diverse and scalable but also fully customizable. Benchmarks should enable full isolation and flexible manipulation of scene properties, such as object density, relative positioning and levels of occlusion levels, through a set of configurable parameters within a large combinatorial space. For example, to evaluate VLM’s robustness to irrelevant objects, the benchmark should be able to systematically increase the number of irrelevant objects, while retaining the task-related objects unchanged. This fine-grained customizablility will allow us moving beyond the aggregate performance metrics (e.g., average task accuracy [yang2025thinking]), to uncover the specific failure modes of VLMs under distinct spatial conditions.
@@
Existing benchmarks face significant limitations in customizablility, scalability, and semantic richness. Real-world datasets [yin2023lamm, hong20233d, jia2025omnispatial, wu2025spatial, chen2024spatialvlm, jia2025omnispatial], although photorealistic, are costly to scale and lack parameterized control. A better alternative is to use synthetic data. Early methods use 3D rendering engines such as Blender [blender] and IsaacSim [isaacSim] to procedurally generate 3D scenes, but lack realism and fail to capture the real-world functional and aesthetic constraints [wang20233d]. Recent techniques address these limitations by using 3D-aware diffusion models [meng2025lt3sd, liu2024pyramid, go2025videorfsplat, go2025splatflow, engstler2025syncity, li2025worldgrow, xiang2025structured] with text [yang2025prometheus, go2025videorfsplat] or image priors [engstler2025syncity, xia2025scenepainter] for visual richness, but lack semantic annotations to generate QA pairs for reasoning tasks and fail to guarantee physical scene coherence. LLM-based scene generation [ran2025direct, bucher2025respace, sun2025layoutvlm, zhang2025scene] can provide better semantic grounding, but falter on generating more complex scenes with correct layouts, due to LLMs’ inherent limitations in spatial reasoning. Optimization-based procedural frameworks, such as Infinigen [raistrick2024infinigen] and ProcTHOR [deitke2022], similarly fail to generate complex and cluttered scenes, and depend on professional tuning in practical use.
@@
To achieve this, we first developed a new LLM-based agentic framework that wraps an advanced optimizer to generate and refine the 3D scene constraints, according to the given prompt of scene descriptions. The refined scene constraints are then transformed into 3D scenes by a custom optimization engine, which iteratively samples and places objects to satisfy all the semantic and physical rules in the scene constraints. In particular, we replace the conventional method of hierarchical optimization [xu2002constraint, yu2011make, merrell2011interactive] with a more flexible cluster-based optimization strategy, to allow creating significantly more complex and dense scenes that were previously intractable. In this way, InfiniBench allows broad but effortless customization of the generated 3D scene from multiple perspectives (e.g., number and type of objects, spatial relationships, room layout, etc), without requiring professional knowledge or complicated tuning.
Another gap between 3D scene generation and VLM benchmarking is that outputs of current scene generation methods, typically 3D models in Blender files or point clouds, are not directly consumable by VLMs. To generate image sequences or video streams that VLMs can properly perceive, we developed a new method of camera trajectory optimization, ensuring that all task-relevant objects and their spatial relationship are clearly visible and covered in camera views. Following the principle of frontier-based exploration [yamauchi1998frontier, lluvia2021active, yamauchi1997frontier], our algorithm incrementally builds the camera path by addressing each task-relevant object in sequence. For each object, it samples and evaluate multiple viewpoints and selects the best one that offers an unobstructed view after passing an occlusion check. A navigable path is then planned to link these viewpoints.
@@
We further showed that customizable benchmarks generated by InfiniBench facilitate in-depth analysis of VLM’s failure modes in various spatial reasoning tasks, including measurement (e.g. metric distance estimation), perspective-taking (e.g. object counting) and spatiotemporal tracking (e.g. identification of order identification) [brown2025sims]. Such customizability in benchmarking would also be a key enabler of enhancing VLM’s spatial reasoning capabilities, via spatially grounded training or prompt tuning.
@@ 2.1 Variability of Scene Complexity
Existing spatial reasoning benchmarks [chen2024spatialvlm, wu2025spatial, jia2025omnispatial] often categorize their data samples merely by task type or define scene complexity in simple terms, such as the number of rooms [chen2025gleam]. Such coarse-grained variability conflates multiple dimensions of the scene complexity, and hence obscures the precise reasons of VLM failures in reasoning tasks, which may include an excessive number of distractors (compositional complexity), an unusually intricate object arrangement (relational complexity), or heavily occluded viewpoints (observational complexity). Instead, in InfiniBench, we allowed much more fine-grained variability of scene complexity in these dimensions, by using separate sets of scene parameters (specified by users in natural language) to explicitly represent complexity levels of these dimensions and generating scenes accordingly.
@@ 2.2 Layout Generation in 3D Scene Generation
The main focus of InfiniBench design is the creation of scene layout, which defines the placement, orientation, and relationships of objects in the scene. One can leverage the reasoning capabilities of LLMs to directly produce a layout and then instruct a 3D engine to place assets in the scene [ran2025direct, bucher2025respace, sun2025layoutvlm, zhang2025scene]. To improve the coherence of generated layouts, some works guide LLMs with structural templates [feng2023layoutgpt] or represent scenes as graphs to facilitate compositional generation [ccelen2024design, deng2025global, gao2024graphdreamer]. However, the generated scene’s complexity is constrained by current LLMs’ flawed spatial logic. As shown in Figure 1, as the number of objects and the intricacy of constraints increase, these methods frequently produce physically implausible layouts, such as those with unrealistic object orientations, out-of-boundary objects or overlapped objects.
To address this challenge and ensure the scalability of generating complex scenes, we proposed a more robust approach that separates high-level planning from low-level execution. Instead of tasking the LLM to generate the exact scene layout, we use LLM to generate high-level constraints about object arrangement. These constraints are then fed into a dedicated optimizer that generates a physically coherent 3D scene. By using optimization-based procedural generation frameworks [raistrick2024infinigen, deitke2022] that are effective in 3D scene synthesis [zhao2021luminous, yu2011make, yu2015clutterpalette], we designed an improved optimizer for generating complex scenes and wrap this optimizer within an LLM-based agentic framework, making the entire pipeline both powerful and user-friendly.
@@ 2.3 Camera Trajectory Generation
Optimization-based methods sample numerous camera parameters with respect to a predefined objective, to maximize the object visibility [bares2000virtual, halper2001camera]. While offering fine-grained control, this approach is expensive and struggles with defining an objective function that robustly captures the visual requirements for complex spatial reasoning tasks. Another alternative is to train a generative AI model [courant2024exceptional, jiang2021camera, wang2024motionctrl]. However, existing datasets for such training are primarily for aesthetically pleasing and object-centric photo shots, e.g., smooth orbits around a single object, rather than trajectories with controlled coverage of multiple objects. Some other methods [xie2023gait, dehghanian2025camera] used reinforcement learning to train an agent for scene navigation without large training data, but needs to retrain the agent for each new scene.
Instead, our approach is inspired by frontier-based exploration [yamauchi1997frontier, yamauchi1998frontier, lluvia2021active], a lightweight technique widely used in robotic navigation and mapping. In InfiniBench, we redefine frontier as object clusters, and sample viewpoints targeting these frontiers accordingly to strategically manage factors such as object occlusion.
@@ 3.1 Agentic Generation of Scene Constraints
A primary barrier to traditional procedural generation of 3D scenes is the need of manually scripting complex scene constraints, a process that requires significant technical expertise, especially when the scene complexity grows. InfiniBench democratizes this process, as shown in Figure 3, by using a LLM agent to translate scene descriptions in natural language into procedural constraints. To ensure that the generated constraints are machine-readable, we teach the LLM agent with domain knowledge of 1) a comprehensive list of available procedural APIs and 2) a set of few-shot examples. The information about procedural APIs covers detailed syntax and functionality with respect to the procedural 3D generation framework being used (e.g., Infinigen [raistrick2024infinigen] and ProcTHOR [deitke2022]). The few-shot examples demonstrate the translation from high-level user requests (e.g., “a cluttered office desk”) to constraint definitions (e.g., set_object_count(monitor, 3), on_top_of(keyboard, desk)). More details about such in-context learning are in Appendix A.
@@ 3.2 Layout Optimization for Complex Scenes
When the scene complexity grows, hierarchical optimization methods used in the existing procedural generation frameworks such as Infinigen [raistrick2024infinigen] and ProcTHOR [deitke2022] would fail, because they first fix the positions of large objects (e.g. desks and beds) and usually result in unsolvable states where there is no space for smaller objects (e.g. chairs around the desk), as shown in Figure 5(a).
@@ 3.3 Camera Trajectory Optimization
Given a generated 3D scene and a spatial reasoning task, the objective of our camera trajectory optimization in InfiniBench is to generate the shortest possible camera path that provides a clear, full and unobstructed view of every task-relevant object. Our method is inspired by frontier-based exploration [yamauchi1997frontier, yamauchi1998frontier, lluvia2021active], a classical robotic navigation technique. Unlike traditional approaches that explore frontiers using occupancy grids, we define “frontiers” as the set of unvisited target objects. We then develop a viewpoint sampling method to find optimal views for these targets while avoiding occlusion. As shown in Figure 6, this process consists of the following steps:
@@
    Path planning: A collision-free path is computed from the current camera position to the selected viewpoint using Dijkstra’s algorithm on a 2D floor plan.
@@
I-Design [ccelen2024design] 0.95 28.4 0.72 1.6 0.4 0.92 27.9 0.68 3.4 5.1 0.90 27.1 0.61 6.9 10.3
Holodeck [yang2024holodeck] 0.97 30.1 0.81 1.9 2.1 0.93 29.5 0.77 4.6 3.1 0.88 28.8 0.71 7.7 9.4
LayoutGPT [feng2023layoutgpt] 0.96 29.8 0.79 1.3 2.4 0.95 29.1 0.74 3.7 4.6 0.93 28.3 0.67 4.5 13.5
Luminous [zhao2021luminous] 0.87 27.5 0.75 0.0 0.0 0.84 26.8 0.70 0.0 0.1 0.42 26.2 0.63 0.0 0.0
Infinigen [raistrick2024infinigen] 0.92 31.0 0.88 0.0 0.1 0.87 30.4 0.84 0.0 0.2 0.64 29.7 0.79 0.0 0.2
@@
I-Design [ccelen2024design] 0.90 28.6 0.73 0.1 0.4 0.87 28.0 0.67 1.3 3.1 0.79 27.2 0.59 5.3 6.4
Holodeck [yang2024holodeck] 0.78 30.4 0.82 0.3 0.2 0.80 29.7 0.76 0.9 5.4 0.75 28.9 0.69 3.1 8.7
LayoutGPT [feng2023layoutgpt] 0.81 29.9 0.80 0.9 0.2 0.83 29.3 0.73 1.7 6.8 0.85 28.5 0.65 5.6 9.6
Luminous [zhao2021luminous] 0.74 27.8 0.76 0.0 0.0 0.64 27.1 0.69 0.0 0.0 0.43 26.4 0.61 0.0 0.0
Infinigen [raistrick2024infinigen] 0.83 31.2 0.89 0.0 0.0 0.71 30.6 0.85 0.2 0.1 0.49 29.9 0.78 0.1 0.2